# Dependence of the critical temperature and disorder in holographic superconductors on superfluid density

### v3

@@ 4 Linear dependence of $T_{c}$ on the superfluid density
In Ref. [15], the authors reveal the dependence of critical temperature ($T_{c}$) in overdoped copper oxide superconductors, specifically La2-xSrxCuO4, on their superfluid density. Using atomic layer molecular beam epitaxy, homogeneous single-crystal films with varying doping levels were synthesized to thoroughly investigate the entire overdoped region. The team measured magnetic penetration depth and zero temperature phase stiffness with high precision across thousands of samples. Results showed a consistent linear decrease of phase stiffness with temperature at all doping levels. In this section, we compute the AC conductivity in our model to investigate the relation between $T_{c}$ and the zero-temperature superfluid density $n_{s}(0)$. The exact AC conductivity in the dilatonic black hole, in which corresponds to the normal phase, was studied in Ref. [49]. The superfluid density $n_{s}$ can be read from the AC conductivity in the superconducting phase. We numerically compute the AC conductivity in the superconducting phase. As a result, we find the linear relation between $T_{c}$ and $n_{s}(0)$ in the $O_{2}$ theory. We also provide the approximate formula for the AC conductivity in the superconducting phase in our model.
@@
where $G(z)$ is a regular function at the horizon $z=1$. According to the prescription of the AdS/CFT correspondence with finite temperatures [55], we can compute the AC conductivity by
@@ 4.1 AC conductivity in the normal phase
We now offer a concise analysis by deriving the exact mathematical formulas for AC conductivity within the normal phase. The normal phase is described by $\Phi(r)=0$, then Eq. (36) becomes
@@
Note that this solution is exactly same as those obtained in Ref. [49] for the 3-charge black hole. The form of the expression can be exchanged by using the property of the hypergeometric function. Using Eq. (39), the AC conductivity is obtained as
@@
The result agrees with the DC conductivity obtained in Refs. [45, 46] with vanishing chemical potential, i.e., in the neutral limit. To obtain the $\omega$-dependent AC conductivity and the $T$-dependent DC conductivity, the nonzero dilaton and the coupling between the dilaton and the Maxwell field are important [49]. If the theory has the S-duality, the conductivity becomes constant as those in the bald black hole that we will show next. For more details about this point, see Ref. [56] and section 3.4.6 of Ref. [57]. We can say that the presence of the non-zero dilaton plays a significant role for the emergence of the linear-$T$ resistivity here.
@@
If $T_{c}$ is larger than $k/(2\sqrt{2}\pi)$, the middle regime disappears. $\rho_{\text{DC}}$ does not jump at $T=k/(2\sqrt{2}\pi)$ because Eq. (44) becomes $1$ there. The behavior of Eq. (46) actually corresponds to the result in the full analysis with backreactions [45].
4.2 Numerical results for the AC conductivity in the SC phase
@@
Figure 6 shows the AC conductivity in the $O_{1}$ theory in the dilatonic black hole. Note that the true ground state is the bald black hole sice $T>k/(2\sqrt{2}\pi)$. We show this plot for the purpose of showing how the AC conductivity behaves if the dilatonic black hole is favored. For the purpose of comparing with the approximation we will explore later, we have scaled $\omega$ by $\expectationvalue{O_{1}}$ here. At the low temperature, the real part of the conductivity exhibits the gap-like structure. This behavior is similar to those in the original model of the holographic superconductor.
Another interesting observable in the superconducting phase is the superfluid density $n_{s}$, which is also called phase stiffness. This quantity can also be understood as the order parameter of the superconducting phase. It can be read from the low-frequency expansion of the AC conductivity as555 $n_{s}$ defined in (47) may be called as the phase stiffness rather than the superfluid density in some studies. In fact, the dimension of $n_{s}$ is $1$, which is different from the dimension of the number density. The number density of the super-fluid components, $\rho_{s}$, might be given by $n_{s}=\rho_{s}/m^{*}$, where $m^{*}$ denotes the effective mass of the charged carrier. In the holographic superconductor models, one may take $m^{*}=\mu$. See, e.g., section 6.4.1 of Ref. [57] for more details.
@@
[…]
Refer to captionRefer to caption
@@
$n_{s}(0)\sim 0.159T_{c}+0.00778\mu,$ (48)
which is shown as the dashed line in Fig. 8. The similar scaling relation is known as the Uemura relation for underdoped materials [58, 59]. It was also observed in the overdoped side of copper oxides [15]. On the other hand, there is a universal scaling law between $n_{s}(0)$ and $T_{c}$ known as Homes’ law given by $n_{s}(0)\propto\sigma_{\text{DC}}(T_{c})\times T_{c}$ in unconventional superconductors [60, 61]. Our result shows the linear relation between $n_{s}(0)$ and $T_{c}$ but $\sigma_{\text{DC}}$ given by Eq. (44) must enter here. Hence, the Homes’ law is not held. We should note we are studying the probe limit. The test of the Homes’ law in the fully backreacted setup in the Gubser-Rocha model was studied in [46].
Regarding Fig. 8, several significant points merit attention. First, it is crucial to note that the precise zero-temperature limit in our geometry presents a subtle issue because the horizon and singularity converge at a single point. To estimate $n_{s}(0)$, we rely on numerical AC conductivity data for a near-zero temperature scenario with $P/r_{h}=0.99999$ because $P/r_{h}=1$ corresponds to $T=0$ from Eq. (6). The empirical values are obtained from solutions where the temperature is approximately an order of $T/T_{c}\sim 10^{-3}$.
@@
4.3 Approximation for the AC conductivity in the SC phase
In this section, we delve into the derivation of an approximate expression for the AC conductivity within the superconducting phase that utilizes the dilatonic black hole geometry as given by Eq.(5). For this purpose, we employ the method developed in Ref. [25].777 We also present another approximate formula in Appendix B.
@@ 4.3 Approximation for the AC conductivity in the SC phase
We focus on the $O_{1}$-theory. In this case, $\Phi$ is written as
@@
$B_{x}^{\prime\prime}(r)+\frac{f^{\prime}(r)}{f(r)}B_{x}^{\prime}(r)+\frac{\omega^{2}}{f(r)^{2}}B_{x}(r)=\left[\frac{2q^{2}e^{-\frac{\phi(r)}{\sqrt{3}}}}{f(r)}\Phi(r)^{2}+\frac{1}{2\sqrt{3}}\left(\frac{f^{\prime}(r)}{f(r)}\phi^{\prime}(r)+\frac{\phi^{\prime}(r)^{2}}{2\sqrt{3}}+\phi^{\prime\prime}(r)\right)\right]B_{x}(r).$ (50)
@@
To solve (52), we utilize the approximate method developed in [25]. Now, we consider replacing $V(r)$ in Eq. (52) with a ‘mean value’ denoted by $\bar{V}$ which is a $r$-constant. Assuming $\omega^{2}>\bar{V}$, we obtain the infalling wave solution as
@@
The integrals diverge due to the infinite volume of the integral region. It will be regularized by introducing a large cutoff.888 In [25], they considered it can be regularized by inserting a small imaginary part of $\omega$. In this study, however, we just consider only the leading contribution at $r=\infty$:
@@
According to the standard prescription of the AdS/CFT correspondence [55], and the Kubo formula, the conductivity is obtained as
@@
If $P=0$, the above result reduces to the approximate formula in the original model of the holographic superconductor, which is given by [26, 25]

### v4

@@ 4 Linear dependence of $T_{c}$ on the superfluid density
In Ref. copperoxides , the authors experimentally reveal the dependence of critical temperature ($T_{c}$) in overdoped copper oxide superconductors, specifically La2-xSrxCuO4, on their superfluid density. Using atomic layer molecular beam epitaxy, homogeneous single-crystal films with varying doping levels were synthesized to thoroughly investigate the entire overdoped region. The team measured magnetic penetration depth and zero temperature phase stiffness with high precision across thousands of samples. Results showed a consistent linear decrease of phase stiffness with temperature at all doping levels. In this section, we compute the AC conductivity in our model to investigate the relation between $T_{c}$ and the zero-temperature superfluid density $n_{s}(0)$. The exact AC conductivity in the dilatonic black hole, in which corresponds to the normal phase, was studied in Ref. Ren:2021rhx . The superfluid density $n_{s}$ can be read from the AC conductivity in the superconducting phase. We numerically compute the AC conductivity in the superconducting phase.
@@
where $G(z)$ is a regular function at the horizon $z=1$. According to the prescription of the AdS/CFT correspondence with finite temperatures Son:2002sd , we can compute the AC conductivity by
@@ 4.1 AC conductivity in the normal phase
We now briefly review the derivation of the corresponding analytic AC conductivity in the normal phase. The normal phase is described by $\Phi(r)=0$, then Eq. (36) becomes
@@
Note that this solution is exactly same as those obtained in Ref. Ren:2021rhx for the 3-charge black hole. The form of the expression can be exchanged by using the property of the hypergeometric function. Using Eq. (39), the AC conductivity is obtained as
@@
The result agrees with the DC conductivity obtained in Refs. Jeong:2018tua ; Homes' law with vanishing chemical potential, i.e., in the neutral limit. To obtain the $\omega$-dependent AC conductivity and the $T$-dependent DC conductivity, the nonzero dilaton and the coupling between the dilaton and the Maxwell field are important Ren:2021rhx . If the theory has the S-duality, the conductivity becomes constant as those in the bald black hole that we will show next. For more details about this point, see Ref. Herzog:2007ij and section 3.4.6 of Ref. Hartnoll:2016apf . We can say that the presence of the non-zero dilaton plays a significant role for the emergence of the linear-$T$ resistivity here.
@@
If $T_{c}$ is larger than $k/(2\sqrt{2}\pi)$, the middle regime disappears. $\rho_{\text{DC}}$ does not jump at $T=k/(2\sqrt{2}\pi)$ because Eq. (44) becomes $1$ there. The behavior of Eq. (46) is qualitatively same as the result in the fully-backreacted analysis Jeong:2018tua , although $\mu$ dependence is dropped out due to the probe limit. It implies that the probe limit analysis still can capture several significant property of this model, such as the linear-$T$ resistivity.
4.2 AC conductivity in the SC phase and the superfluid density
@@
Another interesting observable in the superconducting phase is a superfluid density $n_{s}$. It can be read from the low-frequency expansion of the AC conductivity as
@@
More precisely, $n_{s}$ is proportional to the superfluid density.555 $n_{s}$ defined in (47) may be called as the phase stiffness rather than the superfluid density in some studies. In fact, the dimension of $n_{s}$ is $1$, which is different from the dimension of the number density in $(2+1)$d system. The number density of the super-fluid components, $\rho_{s}$, might be given by $n_{s}=\rho_{s}/m^{*}$, where $m^{*}$ denotes the effective mass of the charged carrier. In the holographic superconductor models, one may take $m^{*}=\mu$. See, e.g., section 6.4.1 of Ref. Hartnoll:2016apf for more details. Remark that the real and the imaginary parts of the conductivity are related with each other by the demand of the causality. Figure 6 shows the numerical results of $n_{s}$ as functions of $T$ for various $k$. The behavior of $n_{s}$ is similar to $\langle O_{i}\rangle$ at low temperatures, see Fig. 4. Unlike Fig. 4, we can see from the numerical results that $n_{s}(T)$ has linear behavior near $T=T_{c}$.
@@
$T_{c}\sim 0.159\,n_{s}(0)+0.00778\,\mu,$ (48)
which is shown as the dashed line in Fig. 7. A similar relation has been also observed in the overdoped side of a copper-oxide film copperoxides . On the other hand, there is a universal scaling law between $n_{s}(0)$ and $T_{c}$ known as Homes’ law given by $n_{s}(0)\propto\sigma_{\text{DC}}(T_{c})\times T_{c}$ in unconventional superconductors Homes:2004 ; Homes:2005 . Our result shows the linear relation between $n_{s}(0)$ and $T_{c}$ but $\sigma_{\text{DC}}$ given by Eq. (44) must enter here. Hence, the Homes’ law is not held. This is stirring result that may imply the $O_{2}$ theory with $\tau=0$ of our model can describe LSCO thin films.
We have to remark some points on Fig. 7. First, it is crucial to note that the precise zero-temperature limit in our geometry presents a subtle issue because the locations of the horizon and the singularity converge at a single point. To estimate $n_{s}(0)$, we rely on numerical AC conductivity data for a near-zero temperature scenario with $P/r_{h}=0.99999$ because $P/r_{h}=1$ corresponds to $T=0$ from Eq. (6). The empirical values are obtained from solutions where the temperature is approximately an order of $T/T_{c}\sim 10^{-3}$.
@@
5 Discussion and Conclusion
To unravel the intricate scaling relations in holographic superconductors, we investigate the relation among the momentum dissipation, the critical temperature, and the zero-temperature phase stiffness in the Gubser-Rocha model within the probe limit framework. This approach has enabled us to examine various fundamental aspects of the holographic superconductor model. We have shown the several properties of the holographic superconductor model, e.g., the critical temperature $T_{c}$, the condensations $\expectationvalue{O_{i}}$, and also the AC conductivity $\sigma(\omega)$ in both phases. In the analysis of the critical temperature, we have found the scaling relation (32) between $T_{c}$ and $k$ for the $O_{1}$ theory. We have also studied the relation between $T_{c}$ and the superfluid density $n_{s}$ at zero temperature in the $O_{2}$ theory as shown in Fig. 7. We found the roughly linear relation, which can be fitted well as Eq. (48) for $\tau=0$. Similar relations have been observed in real experiments, as we will discuss below.
A similar linear relation between $T_{c}$ and $n_{s}(0)$ has also been observed in a recent experiment in a cooper-oxide film copperoxides . Our result may imply that the $O_{2}$ theory with $\tau=0$ of our model can describe such cuprates. Note that it differs from the direct proportionality in underdoped materials, known as Uemura relation Uemura:1989 ; Uemura:1991 . On the other hand, the Homes’ law Homes:2004 ; Homes:2005 is known as the universal scaling relation, which works regardless of whether overdoped or underdoped materials. This relation is expressed as $n_{s}(0)\propto\sigma_{\text{DC}}(T_{c})T_{c}$. The Homes’ law in the holographic superconductor model was tested in Ref. Homes' law in the analysis with backreactions. They concluded that the Homes’ law holds in the $O_{2}$ theory with sufficiently large $\tau$ in this model. In our computation of the superfluid density, we focus on $\tau=0$, where they claim the Homes’ law does not hold in this model. We also remark that the material in copperoxides also does not obey the Homes’ law. It was discussed in dordevic . If we attempt to check the Homes’ law in this model under the probe limit, two problems arise. First, the Homes’ law involves the normal phase DC conductivity at $T=T_{c}$, which is inaccurate in the probe limit. Second, the superfluid density does not converge at low temperatures for $\tau>\tau_{c}$, like in the $O_{1}$ theory, so we cannot obtain $n_{s}(0)$. Even though the Homes’ law does not cover all superconductors, it still valid for various conventional and unconventional superconductors. On the other hand, only the specific holographic model admits the Homes’ law in the limited situations. This mystery of the discrepancy still need to be understood in the future studies.
Our primary emphasis lies on the superfluid density; however, the behavior of the normal fluid density plays a vital role in high-temperature superconductors. Recent studies Gouteraux:2019kuy ; Gouteraux:2020asq have uncovered that in certain classes of holographic superconductors, the normal fluid density persists even down to absolute zero temperature — a phenomenon starkly contrasting the predictions of BCS theory, wherein the normal fluid density vanishes at zero temperature. According to Gouteraux:2020asq , this unconventional persistence occurs when the dynamical exponent $z$ exceeds $d+2$, where $d+1$ represents the spacetime dimension of the boundary theory. Given that the Gubser-Rocha black hole corresponds to $z\rightarrow\infty$, our model likely falls within this category.
Within our setup, the strength of momentum dissipation (or disorder) represented by $k$ introduces an additional parameter to the system compared to the standard holographic superconductor model (as reviewed in bm ). While the role of $k$ is clear in breaking the translational symmetry explicitly in a full analysis with backreactions, its direct experimental correspondence remains uncertain. Typically, translational symmetry is indeed broken in real-world materials, but the intensity of this breakage can vary with changes in other parameters. In experiments involving cuprate superconductors, for instance, hole doping $p$ significantly impacts the material properties. Cuprates exhibit superconductivity within a specific doping range, leading to a characteristic dome-shaped phase diagram along the $p$-axis. It is conjectured that changes in doping levels also lead to variations in $k$. Another factor that can induce momentum dissipation is the strength of defects, which can be manipulated through processes like irradiation. It’s worth noting that translational invariance can also be broken by considering a spatially varying chemical potential $\mu(x)$ Horowitz:2012ky . Works such as Arean:2013mta ; Arean:2015sqa have introduced charged disorder via a random profile of $\mu(x)$ within the probe limit. The presence of $k$ can be viewed as an analogy to these mechanisms, albeit independent of charge. Establishing a precise connection between $k$ and experimentally controllable parameters would be both intriguing and significant.
It is important to acknowledge the limitations imposed by adopting the probe limit in our analysis. Firstly, even when setting $k=0$, the conductivity in the normal phase does not vanish, contrary to what one might expect. This anomaly arises due to the probe limit itself, which inherently leads to a broken translational symmetry. The explicit dependence on the charge density is also dropped out from the conductivity. Secondly, the condensation and the superfluid density sometimes do not converge at low temperatures. Indeed, we encountered a challenge in obtaining solutions for the superconducting phase below the temperature threshold of $T=k/(2\sqrt{2}\pi)$ within the $O_{1}$ theory. It also happens in the $O_{2}$ theory, when we set large $\tau$. In Appendix A, we have numerically verified that the phase boundary calculated under the probe limit does indeed align with the full result when considering a large value of $q$ along with appropriate rescaling. While the probe limit method has the above limitations, it simultaneously simplifies the model’s analysis considerably, making it a practical starting point. Future work, encompassing a more comprehensive treatment of backreactions, is necessary to overcome these constraints and achieve a more accurate depiction of the system’s behavior, especially at low temperatures and for the $O_{1}$ theory.
In this paper, we provided the first comprehensive analysis of the probe limit in this model, as far as we know. Note that the AC conductivity calculation with the black hole solution (5) has been examined in Ren:2021rhx , and the same neutral geometry with $k=0$ was briefly mentioned in Appendix B of Ren:2022qkr . Expanding upon this groundwork, it would be intriguing to tackle more complex scenarios that are challenging to study in a full backreaction analysis, such as spatially inhomogeneous solutions Keranen:2009vi or the temporal evolution of the system Zeng:2016gqj ; Yang:2023dvk . The inclusion of a running dilaton can yield distinct outcomes from standard holographic superconductor setups in the SAdS4 spacetime, although it is essential to remember that the dilatonic black hole becomes the ground state only when the temperature satisfies $T<k/(2\sqrt{2}\pi)$.
Moreover, exploring the impact of magnetic fields in our model promises to be equally fascinating. A dyonic black hole geometry in the Gubser-Rocha model serves as the normal phase vacuum, as recently investigated in Ref. Ge:2023yom . However, in the probe limit, our neutral black hole geometry remains applicable since the Maxwell fields are uncoupled from gravity, as evidenced in analogous studies within the SAdS4 context Ge:2010aa ; Domenech:2010nf . Here, the effect of an external magnetic field solely permeates the system through the coupled Maxwell-scalar sector under the probe limit conditions. By introducing such a magnetic field, it is anticipated that the phase boundaries and condensates will be altered accordingly. Of particular significance is the role played by the dynamical $U(1)$ gauge field in capturing various magnetic field effects Domenech:2010nf ; Salvio:2012at ; Salvio:2013jia . Holographic superconductors featuring dynamical $U(1)$ gauge fields are known to emulate type II superconductors, and the penetration depth computed in their presence constitutes a crucial quantity that can be more directly compared to experimental findings. These areas remain ripe for future research endeavors.
Acknowledgments
Appendix A Checking the correctness of the probe limit
In this section, we compare the $T_{c}$$k$ curves obtained by the probe limit analysis with those obtained for general $q$. Firstly, we briefly review the charged black hole solutions of our model (1), and its thermodynamic stability. From the full action (1), the equations of motion are obtained as
$D_{\mu}D^{\mu}\Phi-B(\phi)\Phi=0,\quad\nabla^{2}{\psi_{I}}=0.$ (51)
The Einstein’s equation is
The normal phase solution is obtained as the following charged dilatonic black hole solution Jeong:2018tua :
$ds^{2}=-f(r)dt+\frac{dr^{2}}{f(r)}+h(r)(dx^{2}+dy^{2}),$ (53)
with
where $r_{h}$ is the location of the black hole horizon and $P$ is a physical parameter. Unlike Eq. (5), $r_{h}$ and $k$ are independent parameters here. Note that this is the same solution as those in Ref. Homes' law , but the coordinate is different. Writing the radial coordinate in Homes' law as $\tilde{r}$, we obtain the relation between ours and their conventions as $\tilde{r}+Q=r$, $\tilde{r}_{h}+Q=r_{h}$ and $Q=P$. The Hawking temperature and the chemical potential are related to the parameters by
One can see that the neutral limit of Eq. (5) is achieved by setting $r_{h}=k/\sqrt{2}$, whereas $P=0$ leads Eq. (7). $P/r_{h}\to 1$ corresponds to the extremal limit. For the thermodynamic analysis, the grand potential (density) of this solution is given by Caldarelli:2016nni ; Kim:2017dgz
The grand potential implicitly depends on $\mu$ and $T$ via $P$ and $r_{h}$.
On the other hand, the model also admit the following solution: Andrade:2013gsa
One can see $\mu=0$ leads the neutral solution of Eq. (7). The grand potential of this solution is given by Andrade:2013gsa
The grand potential implicitly depends on $T$, again.
Let us now consider which solution and parameter range correspond to the ground state of the normal phase. From Eq. (55), there are two distinct parameter regions yielding $T,\mu>0$ for the dilatonic black hole; i) $0<P<r_{h}$ and $0<k<\sqrt{2}r_{h}$ ii) $P<0$ and $\sqrt{2}r_{h}<k<\sqrt{6}r_{h}$.666 In terms of the original parameter $\tilde{r}_{h}$, the ranges of such parameter regions become more complicated. The both choice of the patches cover the entire region of the physical parameter-space like $(k/T,\mu/T)$. On the other hand, $k<\sqrt{6r_{h}^{2}-\mu^{2}/2}$ is obtained from Eq. (58) for the bald black hole. To determine the ground state, we need to compare the grand potential among these cases. It was investigated in Kim:2017dgz , and the answer is that the $P>0$ patch of the dilatonic black hole is always ground state. Figure 8 shows the competition of the grand potential among these solutions for fixed $\mu/T=5.0$. $\Omega_{\text{bald}}$ denotes Eq. (59). $\Omega_{P>0}$ and $\Omega_{P<0}$ denote Eq. (56) in the $P>0$ and $P<0$ patches, respectively. The curves exhibit that $\Omega_{\text{bald}}-\Omega_{P<0}<0$, $\Omega_{P>0}-\Omega_{P<0}<0$ and $\Omega_{P>0}-\Omega_{P<0}<\Omega_{\text{bald}}-\Omega_{P<0}$. It reads $\Omega_{P>0}<\Omega_{\text{bald}}<\Omega_{P<0}$. Therefore, we conclude that the $P>0$ patch is the ground state, which is thermodynamically favored Kim:2017dgz . We have checked this behavior of the grand potentials does not change for another choice of $\mu/T$.
Figures 9 and 10 show the phase boundaries in the $\frac{T}{q\mu}$$\frac{k}{q\mu}$ plane for various $q$ in each theory. The results for $q=\infty$ denote the results obtained by using the neutral solution (5) and (7), i.e., the results in the probe limit. Here, we set $\tau=0$ for simplicity, and show only the relevant results in the ground states. In both cases, we can see that the full results approach those in the probe limit as we increase $q$. These direct comparisons imply that we need considering the phase transition of the black hole in the probe limit.
Figure 9: Superconducting phase boundaries of the $O_{1}$ theory in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for various $q$. The results for $q=\infty$ denote the results in the probe limit. The $q=10$ curves almost overlap with the $q=\infty$ curves. The dotted line shows $T_{c}=k/(2\sqrt{2}\pi)$.
Figure 10: Superconducting phase boundaries of the $O_{2}$ theory in the $\frac{T_{c}}{q\mu}$$\frac{k}{q\mu}$ plane for various $q$. The results for $q=\infty$ denote the results in the probe limit. The result for $q=6$ corresponds to one studied in Ref. Homes' law . The dotted line shows $T_{c}=k/(2\sqrt{2}\pi)$.
Appendix B Approximate formulas of the AC conductivity in the SC phase
In this section, we derive approximate expressions of the AC conductivity within the superconducting phase where the dilatonic black hole geometry (5) is favored. We show two results obtained by using two distinct methods developed in Herzog:2009xv ; Siopsis:2010uq ; Gary .
B.1 A method using the tortoise coordinate
@@
First, we employ the method developed in Ref. Herzog:2009xv ; Siopsis:2010uq ; Gary . We focus on the $O_{1}$-theory. In this case, $\Phi$ is written as
@@
$B_{x}^{\prime\prime}(r)+\frac{f^{\prime}(r)}{f(r)}B_{x}^{\prime}(r)+\frac{\omega^{2}}{f(r)^{2}}B_{x}(r)$ (61)
$=\left[\frac{2q^{2}e^{-\frac{\phi(r)}{\sqrt{3}}}}{f(r)}\Phi(r)^{2}+\frac{1}{2\sqrt{3}}\left(\frac{f^{\prime}(r)}{f(r)}\phi^{\prime}(r)+\frac{\phi^{\prime}(r)^{2}}{2\sqrt{3}}+\phi^{\prime\prime}(r)\right)\right]B_{x}(r).$
@@
To solve (63), we utilize the approximate method developed in Siopsis:2010uq . Now, we consider replacing $V(r)$ in Eq. (63) with a ‘mean value’ denoted by $\bar{V}$ which is a $r$-constant. Assuming $\omega^{2}>\bar{V}$, we obtain the infalling wave solution as
@@
The integrals diverge due to the infinite volume of the integral region. It will be regularized by introducing a large cutoff.777 In Siopsis:2010uq , they considered it can be regularized by inserting a small imaginary part of $\omega$. In this study, however, we just consider only the leading contribution at $r=\infty$:
@@
According to the standard prescription of the AdS/CFT correspondence Son:2002sd , and the Kubo formula, the conductivity is obtained as
@@
If $P=0$, the above result reduces to the approximate formula in the original model of the holographic superconductor, which is given by Gary ; Siopsis:2010uq